# DP-GTR: Differentially Private Prompt Protection via Group Text Rewriting

### v1

@@ A.1 DP-guaranteed Paraphrasing Proof
2:  $\text{LM}\leftarrow\texttt{clipLogits}(u,\mathbf{b})$
3:  $\text{Temp}\leftarrow\bigl{(}\tfrac{2\bigl{(}b_{max}-b_{min}\bigr{)}}{\epsilon}\bigr{)}$
@@
A.2 Epsilon and sensitivity clip
The sensitivity bound is the other critical theoretical parameter. Following prior work, we adopt a pre-clipping strategy following the previous studies Igamberdiev and Habernal (2023). Specifically, we randomly sample 1,000 examples from the CSQA training dataset and perform the DP-Prompt paraphrasing task while recording all logits. We then compute the mean ($\mu$) and standard deviation ($\sigma$), and define the sensitivity bound as $(\mu,\mu+4\sigma)$ to better preserve high-value logits (Meisenbacher et al., 2024b).
@@ A.2 Epsilon and sensitivity clip
Temperature DP-MLM (ROBERTA) DP-Paraphrase (GPT-2) DP-Prompt (T5) Ours-NDP (llama) Ours-JEM (llama)
0.1 390.0 1760.0 534.2 194.0 194.0
0.15 260.0 1173.3 356.1 129.3 129.3
0.2 195.0 880.0 267.1 97.0 97.0
0.25 156.0 704.0 213.7 77.6 77.6
0.5 78.0 352.0 106.8 38.8 38.8
0.75 52.0 234.7 71.2 25.9 25.9
1.0 39.0 176.0 53.4 19.4 19.4
1.25 31.2 140.8 42.7 15.5 15.5
1.5 26.0 117.3 35.6 12.9 12.9
Table 2: Epsilon values for different methods across temperatures

### v2

@@ Appendix A DP-guaranteed Paraphrasing Proof
[…]
@@
Appendix B Epsilon and Sensitivity Setting
The sensitivity bound is the other critical theoretical parameter. Following prior work Igamberdiev and Habernal (2023), we adopt a pre-clipping strategy. Specifically, we randomly sample 1,000 examples from the CSQA training dataset and perform the DP-Prompt paraphrasing task while recording all logits. We then compute the mean ($\mu$) and standard deviation ($\sigma$), and define the sensitivity bound as $(\mu,\mu+4\sigma)$ to better preserve high-value logits (Meisenbacher et al., 2024b). We also detail the exact logits clipping ($b_{min},b_{max}]$) range used in Table˜4.
Table 4: Logits clipping bounds ($b_{min}$ and $b_{max}$) for different DP-based models.
Model $b_{min}$ $b_{max}$
DP-MLM (RoBERTa) -3.2093 16.3048
DP-Paraphrase (GPT-2) -96.8525 -8.7477
DP-Prompt (Flan-T5) -19.2271 7.4832
Ours (Llama-3.1-8B) 22.9443 32.6274
@@ Appendix B Epsilon and Sensitivity Setting
Table 5: Epsilon values for different methods introduced in Section˜5.2 across temperatures (T).
T DP-MLM DP-Paraphrase DP-Prompt Ours
RoBERTa GPT-2 FLAN-T5 Llama
0.10 390.0 1760.0 534.2 194.0
0.15 260.0 1173.3 356.1 129.3
0.20 195.0 880.0 267.1 97.0
0.25 156.0 704.0 213.7 77.6
0.50 78.0 352.0 106.8 38.8
0.75 52.0 234.7 71.2 25.9
1.00 39.0 176.0 53.4 19.4
1.25 31.2 140.8 42.7 15.5
1.50 26.0 117.3 35.6 12.9
Additionally, in the setting of Ours-JEM, $\varepsilon_{2}$ for JointEM is set to 2.
Appendix C Computational Costs
DP-GTR is an inference-side, single-prompt privacy protection approach that demonstrates lower overall computational resource consumption compared to methods requiring model training Hong et al. (2024). The primary source of computational overhead stems from the grouping step involved in paraphrasing. Below, we provide a detailed computational analysis.
  • Computation Efficiency: DP-GTR is designed to support parallel generation of paraphrased texts, which can substantially reduce inference time from $O(n+1)$ to $O(2)$ when there are sufficient computational resources.
  • Memory Efficiency: When parallel processing is not feasible, GTR can operate in serial mode. In this case, only the sentence with the lowest perplexity and the corresponding privacy keyword votes are retained in memory, thereby minimizing memory consumption from $O(n)$ to $O(1)$.
  • User-End Efficiency: For user-end environments with low throughput and limited computational resources, DP-GTR does not require model training and offers particular advantages in both memory and computation efficiency.
While we acknowledge that DP-GTR introduces additional computational overhead relative to baseline inference, we argue that this trade-off is justified. The enhanced privacy protection offered by DP-GTR is particularly important for practical deployments, given the increasing privacy risks associated with large language models.
Appendix D Adversarial Attack
We consider a threat model in which the attacker has full capability at the cloud end at Preliminaries 3. Accordingly, we detail both static and adaptive attack strategies. In the static attack, the adversary directly prompts the LLM to reconstruct the original question using the template shown in Template D. For the adaptive attack, the adversary replicates the entire workflow, including candidate paraphrase generation, private consensus keyword extraction, and reference selection, and then combines this information to recover the original question. The adaptive prompt recovery template is provided in Template D.
Static Prompt Template Given a paraphrased question: $\hat{P}$> Convert to the original question:
Adaptive Prompt Template Refer the paraphrased question to generate the original question: $\hat{P}_{low}$> Use following tokens: $\hat{w}_{1},\hat{w}_{2},...,\hat{w}_{k}$> Generated original question:
Appendix E Explainable Case
To illustrate our approach, we present an explainable case below, which is forked from our open-source repository and generated by DP-GTR-NDP (GPT-3.5). The Document is the original question, followed by 10 DP-guaranteed paraphrases. From these, we extract the Private Keywords and identify the lowest-perplexity Reference, ultimately producing a differentially private, protected document.
DP-GTR Example Document: In which year, john f. kennedy was assassinated? 10 DP-guaranteed Paraphrases: In what year was John F. Kennedy killed? What year was John F. Kennedy assassinated? During which year did John F. Kennedy get assassinated? In what year was John F. Kennedy killed? What year did John F. Kennedy get assassinated? In what year was John F. Kennedy killed? What year did John F. Kennedy get assassinated? In what year was John F. Kennedy killed? What year did John F. Kennedy get shot? In what year did John F. Kennedy get killed?
Private Keywords: What, get, Kennedy, killed, In, John, assassinated, F., year, During Reference: What year was John F. Kennedy assassinated? Final Generated Document: When did the tragic event occur involving the 35th President of the United States?